Solve each quadratic equation by completing the square.
1.X^2+4x+2=0
2.X^2+8x=1
3.X^2+2=6x+4

Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

Solve each quadratic equation by completing the square.
1.X^2+4x+2=0
x^2+4x+? = -2+?
x^2+4x+4 = -2+4
(x+2)^2 = 2
---
x+2 = sqrt(2) or x+2 = -sqrt(2)
x = -2+sqrt(2) or x = -2-sqrt(2)
---------------------------------------
2.X^2+8x=1
x^2+8x+? = -1+?
x^2+8x+16 = -1+16
(x+4)^2 = 15
x+4 = sqrt(15) or x+4 = -sqrt(15)
---
x = -4+sqrt(15) or x = -4-sqrt(15)
-------------------------------------------
3.X^2+2=6x+4
x^2-6x = 2
x^2-6x+? = 2+?
x^2-6x+9 = 2+9
(x-3)^2 = 11
x-3 = sqrt(11) or x-3 = -sqrt(11)
x = 3+sqrt(11) or x = 3-sqrt(11)
==================================
Cheers,
Stan H.
============

Answer by nerdybill(7384)

1.X^2+4x+2=0
X^2+4x = -2
X^2+4x+4 = -2+4
(x+2)(x+2) = 2
(x+2)^2 = 2
(x+2) = sqrt(2)
x = sqrt(2) - 2
.
2.X^2+8x=1
X^2+8x+16 = 1+16
(x+4)(x+4) = 17
(x+4)^2 = 17
(x+4) = sqrt(17)
x = sqrt(17) - 4
.
3.X^2+2=6x+4
X^2=6x+2
X^2-6x = 2
X^2-6x+9 = 2+9
(x-3)(x-3) = 11
(x-3)^2 = 11
(x-3) = sqrt(11)
x = sqrt(11) + 3